OpenStudy (anonymous):

which equation has a graph perpendicular to the graph of 7x = 14y - 8 ?

OpenStudy (anonymous):

do you have choices for the question?

OpenStudy (anonymous):

A) y = -2x - 7 B) y = - 1/2x + 1 C) y = 1/2x + 1 D) y = 2x + 9

OpenStudy (anonymous):

set the above equation in to y=mx+b.. m being the slope, and to find the perpendicular, you just need to find the reciprocal of the slope with the opposite sign

OpenStudy (anonymous):